Run Group C

From clas12-run
Jump to: navigation, search
[edit]

NOTE: All NON-JLab numbers MUST be dialed with (757) area code. When calling from a CH landline, dial "9" first. To call from outside, all 4-digit numbers must be preceded by 757-269

Person Phone Number
Run Coordinator 9-757 575-7540 (cell)
Physics Division Liaison 9-757 876-1789 (cell)
Shift Expert 5244 or 9-757 329-4846 (cell)
Shift Worker 5126
MCC-OPS 7048
Crew Chief 7045
Crew Chief 9-757 876-3367 (cell)
Program Deputy 9-757 876-7997 (cell)
Engineering on-call 9-757 748-5048 (cell)
Engineering on-call secondary 9-757 897-9060 (cell)
Radcon (needed for target cell swaps) 9-757 876-1743 (cell)
Gate House Guard 5822
Location Phone Number
Hall-B Floor 5165
Hall-B Space Frame 5170, 5171
Hall-B Forward Carriage 5371
Hall-B Counting Room 5245 (Shift1), 5126 (Shift2)
Hall-B Counting Room 5244 (Expert)
Hall-B Gas Shed 7115
System Phone Number On-Call Person
DC (757) 748-5048 (cell) Engineering On-Call
SVT + MM (757) 541-7539 (cell) Yuri Gotra Primary
SVT + MM (757) 753-7769 (cell) Rafo Paremuzyan (secondary)
ECAL (757) 810-1489 (cell) Cole Smith
FTOF/CTOF (757) 344-7204 (cell) Jose Carvajal
CND (757) 344-7204 (cell) Jose Carvajal
HTCC-LTCC (757) 344-7174 (cell) Youri Sharabian
FT (757) 344-1848 (cell) Raffaella De Vita
BAND (757) 310-7198 (cell) Florian Hauenstein
RICH (757) 344-3235 (cell) Valery Kubarovsky
DAQ (757) 232-6221 (cell) Sergey Boiarinov
Slow Controls (757) 748-6922 (cell) Nathan Baltzell
Beamline (757) 303-3996 (cell) Eugene Pasyuk
Pol. Target On-Call (757) 218 2266 (cell) James Maxwell
Polarized Target (757) 871-5374 (cell) James Brock
Target (not used) (757) 822-9586 (cell) Bob Miller Primary
Target (not used) (757) 897-9060 (cell) Denny Insley Secondary
Counting House C x6000
Counting House D x5504

In case of a problem with superconducting magnets call Engineering on-call first, particularly if it is related to cryo.
For all other superconducting magnet problems call magnet group.
Their schedule and phone numbers are posted on the wall next to the Saclay target computer.




Click Here to edit Phone Numbers. Note, you then also have to edit the current page to force a refresh.



CLAS12 Run Group C, Summer 2022

Daily RC Meetings at 14:00: CH Room 200C, and on Zoom

Zoom information: Link, Meeting ID: 160 867 4253, Passcode: 786506

Run Coordinator: Yelena Prok

  • (9-)757 575-7540

PDL: Eugene Pasyuk

  • (9-)757 876-1789 , x6020 office

Run Group Coordinator: Sebastian Kuhn

  • (9-)757 639-6640

Run Coordinator and Target Expert Schedule

Longterm Run Plan

IMPORTANT NOTES:

  • In case of medical or fire emergency, call 9-911 from a LANDLINE phone in the Counting House
  • If you have symptoms or a positive COVID test, do NOT come to your shift - instead, inform the PDL and the RC and JLab medical services immediately so that replacements can be arranged
  • Recently there were several reported cases of COVID on the JLab campus. Everyone is strongly encouraged to wear a face mask while in the counting-house.
  • If needed, worker shifts can also be taken from off-site - again, inform the PDL and RC. Also, make sure that you follow the instructions under the "Worker Shift" Tab - you need some lead time to set up!
  • Before leaving the hall and going to beam permit check that dome lights, forward carriage lights, and pie tower lights are OFF
  • NEW: We have a new policy NOT to leave the beam viewer in the beam at all times, to avoid damage to electronics, cameras etc. Please REMOVE the beam viewer unless we are actively tuning the beam or calibrating beam position/raster.
  • NEW: We have a new FSD condition that the raster is operating.
    • IF there is any ammonia sample (NH3 and ND3) in the target, the raster MUST be operating at all times. The Raster FSD must be UNMASKED. Orbit locks MUST be OFF (except for brief moments to restore beam position).
    • For beam tuning or other non-rastered beam, the Raster FSD must be MASKED.
  • NOTE: It should no longer be necessary to quit RasterMon at the end of each run and restart after a successful "Go" for the next run, but if DAQ crashes, you SHOULD follow these steps, and the same for mon12. Make sure there is ONLY ONE INSTANCE of RasterMon AND mon12 running anywhere in the Counting House.
  • REMINDER FOR ALL EXPERTS: Please take care to fill out the BTA carefully and correctly. The info from EPICS is not always correct, e.g. when taking data with low beam current, you should change the ABU to the proper (estimated) time by hand instead of simply accepting "0" if you are actually taking data.

Run Plan: August 2022

GENERAL INFORMATION

  • We should change the HWP ("lambda/2 plate"), after a full day of useful data on NH3 or ND3. WE MUST INFORM THE OTHER HALLS WHENEVER WE DO THAT
  • We should reverse the target polarization after every TWO days of good data taking. So the sequence is "1 day on ammonia, swap HWP, one day on ammonia, flip target polarization, one day on ammonia, swap HWP, one day on ammonia".
  • Moller runs should be taken with 10 nA.
  • Do a RICH Recovery procedure only once a day during day shift (in between runs) if everything is okay. If a RICH problem occurs, you might have to do the RICH Recovery procedure also (see below).
  • After each change of running conditions (target type, luminosity): Please check that the trigger alarms (on rate/nA) are set appropriately and change them as needed.

ONGOING ISSUES

  • x-Raster power supply oscillations

SHORT RANGE PLAN

Wednesday 10th August

  • Owl: Take production data on ND3 negative at 4 nA HWP OUT
  • Day: Take production data on ND3 negative at 4 nA HWP OUT
    • Change HWP to IN during day shift (inform Halls C and D about an hour beforehand)
  • Swing: Take production data on ND3 negative at 4 nA HWP IN

Thursday 11th August

  • Owl: Take production data on ND3 negative at 4 nA HWP IN
    • 6:30 AM Beam will be off for Machine Development Day
    • Hall B will go to Restricted Access
  • Day: Machine Development Day
    • Change of target material to NH3
  • Swing: Tune beam and adjust the raster parameters to accommodate for the smaller cell.
    • Take production data on NH3 at 4 nA HWP IN

Friday 12th August

  • Owl: Take production data on NH3(+) at 4 nA HWP IN
  • Day: Plan to replace target material with another NH3 sample
    • Optimize raster parameters for the 15 mm cell until 8:30.
    • Ask for Controlled Access at 8:30am. Call RadCon to inform them of the target work.
  • Swing:Take production data on NH3 at 4 nA HWP IN

5-pass (10.55 GeV) Running conditions

  • Torus at 100% current, inbending electrons. Setpoint = +3770 A (Readback = +3766 A)
  • Solenoid setpoint +2417.5 A
  • Establish optimal compromise between raster size (if possible, 7 mm radius), luminosity (goal: <5 nA, nominally 4.0 nA), and background.
  • Run duration 100M events or 4 hours, whichever comes first.
    • For empty target runs, increase beam current as much as possible (up to 15 nA).

DAQ configuration

  • Configuration PROD66, trigger file rgc_300MeV_v1.3_no_DC.trg
  • NOTE: UNLESS a RICH problem crashes the DAQ, do NOT interrupt an otherwise smooth run to do a RICH recovery - instead do this in between runs.
  • NOTE: BOTH the shift expert AND the shift worker MUST check and reset monitoring histograms at least once every 30-60 minutes to make sure there are no major detector problems.Watch the raster pattern!
  • At the end of each run, follow the STANDARD DAQ RESTART SEQUENCE:
    • "end run", if the run ended correctly then: "prestart", "go"
    • if the run did not end correctly or if any ROCs had to be rebooted:
      • Quit ALL running instances of mon12 and rastermon
      • "cancel", "reset"
      • "configure", "download", "prestart", "go"
      • restart mon12 and rastermon
  • After each step, make sure it is completed in the Run Control message window. If a ROC has crashed, find which one it is, issue a roc_reboot command ON JUST THAT ROC and try again. Contact the DAQ expert if there are any questions.

References and Standards

Solenoid set-point of 2417.5 A results in read back 2419.57 A and exactly 5.00 T field on center Raster settings

  • x: Vpp=2.2 V offset =-0.3V (changed Vpp x & y from 2.1 to 2.2 V on July 19 for ND3)
  • y: Vpp=2.2 V offset =+0.25V (changed from +0.4V on July 14, changed from -0.1 on July 1 at 21:00)
  • Fundamental Period = 1 s
  • The correct raster calibration constants should be automatically loaded by rastermon - see https://logbooks.jlab.org/entry/4004055 - if not, enter them through "Config" by hand.

Nominal Beam Positions

  • 2H01, X = -2.3 mm, Y = +1.2mm (Changed from 1.8 mm 7/14/22)
  • 2C24, X = -2.0 mm, Y = +0.6 mm
  • Orbit lock ONLY if raster is OFF

FSD Thresholds (to be updated)

  • Upstream: 2000 Hz
  • Midstream: 2000 Hz
  • Downstream: 100000 Hz
  • BOM: 1000000 Hz
  • 5 ms dwell time

Reference Harp Scans for Beam on Faraday Cup 2H01 [1]

Reference Harp Scans for Beam on Tagger dump 2C21 [2], tagger harp [3]

Reference Monitoring Histograms

Counter rates

  • Upstream counters integrated rates: 0-15 Hz (acceptable up to 100 Hz are acceptable) @50 nA.
  • Midstream counters: 10-20 Hz (acceptable up to 50 Hz) @50 nA.
  • Counting rates of ~ hundreds of Hz may indicate bad beam tune or bleed-through from other Halls.

Known issues

  1. ECAL: ECAL_HV_SEC2_WI_E36 is off as it does not hold voltage
  2. tdcecal6 and tdcecal1 rocs: These rocs could show errors in the DAQ screen. If so, check roc-rates, trigger rates and monitoring plots. If all looks okay, continue data taking.
  3. DC: hole in sector 3, superlayer 6, would likely need a major physical repair on the chamber itself


NOTE: BOTH the shift expert AND the shift worker MUST check monitoring histograms at least once every 30 - 60 minutes to make sure there are no major detector problems.

  • Follow the [Short Term Schedule] as outlined by RC
    • includes Run Plan, DAQ configuration, and References and Standards
  • Keep shift summary up to date in HBLOG. Record all that happens.
  • Polarized target issues: In all cases contact the polarized target expert on call. Examples to keep an eye on:
    • Ice on the run valve
    • NMR running
      • The JLab Polarization Display is seen via a VNC window in the counting room. Here, the NMR Controls' blue bar should be moving, indicating that the NMR is running and the polarization is measured. If the NMR is not running, contact the expert on call.
    • Alarms from System RG-C Polarized Target (alarm handler). If you get an alarm, ...
  • Do a RICH recovery (press the RICH Recovery button on the GUI) once daily during day shift, between runs. Please log in shift summary.
    • Note the automated "RICH Recovery" sequence already confirms RICH4 is up and accepting ssh connections before claiming success. And it also prints out that you have to "Cancel->Reset->Configure->Download->Prestart" before starting the next run.
  • Fill out BTA hourly. Click "Load from EPICS" to automatically fill the left side then correct as needed.
    • Accelerator beam time:
      • ABU: acceptable beam used (eg harp scans, calibrations, trigger settings, running)
        • The info from EPICS is not always correct, e.g. when taking data with low beam current, you should change the ABU to the proper (estimated) time by hand instead of simply accepting "0" if you are actually taking data.
      • BANU: beam acceptable but not used (eg radiator change or our problems)
      • BNA: beam not acceptable (eg tuning, drifting, not focused)
    • Experiment beam time
      • ER: experiment ready
      • PCC: planned configuration change (when beam is masked while radiators are moved, DAQ (roc rebooting, restarting runs)), if more than 5 minutes, put a short comment, eg harp scans or DAQ.
      • UED: unplanned experiment down (hardware failure)


In case of other questions or uncertainties of what to do, call the RC or PDL

(remote) shift worker: see dedicated Worker Shift tab above.

For REMOTE Shifts (offsite):

Shift Worker Tasks

Once per run:

  • Screenshots: submit to logbook only if you capture the information with beam ON
    • Beamline Overview, in VNC or Web Browser
    • CLAS12 Trigger rates and DAQ status
    • RGC Polarized Target GUI - under Beamline on CSS
  • Upload monitoring plots to logbook
    • Standard spectra (mon12): look at the spectra: Reference, current, and 5 most recent runs. You can choose a different reference run, the reference numbers refer to logbook entries.
    • Raster Monitoring: upload RasterMon plots to the logbook (use "Log Entry" button on RasterMon screen at bottom-right corner)
  • Make a log entry with the photograph of the polarized target run valve (hallbcam04)

Once per shift:

  • Fill the shift checklist
  • Log screen shots of RICH temperature and scaler maps. Log in RICH logbook.

Regularly:

  • reset & check spectra frequently (every 30 - 60 minutes)
    • Some plots in the online mon12 have a choice of histogram sets per sector, select buttons at the bottom of the plot. That means you have to look at all sectors for these histograms.
    • The logbook entry shows all histograms.
  • Log the strip charts in time intervals according to the x-axis time span, using the liveplot configuration Beamline4Logbook_visibleH
    • how to get the strip charts: CS-Studio Tools: Strip Charts, livePlot
      • right-click on the livePlot screen and select "Load a Configuration"
      • the configuration "Beamline4Logbook_" provides the beam current, position information and halo counter rates
      • right-click on the screen and select "Make Elog Entry", make sure to log to HBLOG and not ELOG
    • in Web Browser: Upstream/Midstream Halo Counters, Downstream Halo Counters, Beam Currents, BPM Positions

Webcams

Manuals

Epics on the web

Live Monitoring Links



Clas12design.png